# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

### v2

@@ 3.2 Bug Oracles
A mempool execution from initial state $st_{0}$ to end state $st_{n}$ under transaction sequence $ops$ is a successful ADAMS eviction attack, if.f. 1) the attack damage is full, that is, end state $st_{n}$ contains only adversarial transactions and initial state $st_{0}$ contains only benign transactions, and 2) the attack cost is low, that is, the total adversarial transaction fees in the end states are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the victim transaction fees in the initial state. Formally,
$st_{0}\cap{}(st_{n})$ $=$ $\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\epsilon\sum_{tx\in{}st_{0}}tx.fee$ (1)
@@
A mempool execution from initial state $\langle{}st_{0},dc_{0}\rangle{}$ to end state $\langle{}st_{n},dc_{n}\rangle{}$ under transaction sequence $ops$ is a successful ADAMS locking attack if.f. 1) the attack damage is full, that is, the victim normal transactions in $ops$ are completely declined by the mempool storing adversarial transactions, where normal transactions and adversarial transactions are sent by different accounts, and 2) the attack cost is low, that is, the total adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the fees of victim transactions declined. Formally,
$\cup_{txd\in{}dc_{n}}{txd.sender}$$\cap$ $\cup_{tx\in{}s_{n}\cup{}ops\setminus{}dc_{n}}{tx.sender}=\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\lambda\sum_{tx\in{}dc_{n}}tx.fee$ (2)
@@ 4 Stateful Mempool Fuzzing by mpfuzz
4.1 Motivating Observations
Tyr’s test oracle detects property violations on post-consensus states. Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low Gas price can be dropped by Ethereum mempools and will not be eventually executed. Besides, they model double-spending transactions as the only type of invalid transaction. In Ethereum, double-spending transactions are of the same nonce. And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
Challenges: In stateful mempool fuzzing, the essential design problem is how to generate the next input transaction given a current “state”. State $st$ is the collection of transactions in the mempool and those included in the blockchain.
The key challenge is the large input space of transactions. Recall that an Ethereum transaction consists of multiple attributes (sender, nonce, Gas price, value, etc.), each defined in a large domain (e.g., 64-bit string). Exploring the raw transaction space is hard; the comprehensive search is inefficient, and randomly trying transactions as done in the state-of-the-art blockchain fuzzers (in § 2) is ineffective.
4.2 Proposed Fuzzer Design
4.2.1 Transaction Symbolization
Intuition: We propose symbolizing transactions for efficient and effective mempool stateful fuzzing. Our key idea is to map each group of concrete transactions triggering an equivalent mempool behavior into a distinct symbol so that searching one transaction is sufficient to cover all other transactions under the same symbol. Transaction symbolization is expected to reduce the search space from possible concrete transactions to the symbol space.
Table 2: Symbols, transactions and associated costs. Tx refers to the instantiated transaction under a given symbol.
Symbol Description
Tx
$cost()$ $opcost()$
$\mathcal{N}$ Benign
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$
$3$ $3$
$\mathcal{F}$ Future
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{P}$ Parent
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$
$[4,m]$ $[4,m]$
$\mathcal{C}$ Child
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
$m+1$ $1$
$\mathcal{O}$ Overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{L}$ Latent overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{R}$ Replacement
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$
00 00
Specification: Suppose the current state contains adversarial transactions sent from $r$ accounts $A_{1},\dots,A_{r}$. These accounts have an initial balance of $m$ Ether, where $m$ equals the mempool capacity (say $m=1000$).
@@ 4.2.1 Transaction Symbolization
Symbol $\mathcal{F}$ defines the transaction subspace of any future transaction sent from an adversarial account; the future transaction is defined w.r.t. the current state. mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, Gas price $m+1$ wei, and any adversarial account $*A*$. The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$.
Symbol $\mathcal{P}$ defines the transaction subspace of any parent transaction from an adversarial account w.r.t. the current state. mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$. Here, the transaction is fixed with a Gas price in the range $[4,m]$ wei, value at $1$ Ether, and nonce at $1$.
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. The instantiated child transaction is fixed at Gas price of $m+1$
Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. Recall that all adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{L}$ defines the transaction subspace of any adversarial latent-overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. All adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{R}$ defines the transaction subspace of any adversarial “replacement” transaction w.r.t., the state. Unlike the symbols above, the transactions under Symbol $\mathcal{R}$ must share the same sender and nonce with an existing transaction in the current state $st$. mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$. Here, we simplify the problem and only consider replacing the transaction with nonce $1$.
@@
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 1, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$).
@@ 4.2.2 State Search Algorithm
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [16]). A seed consists of a symbolic input $ops$, which consists of transaction symbols, and a symbolic state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
@@
Mutation feedback: We describe how mempool states are symbolized before presenting state-based feedback. In mpfuzz, a symbolized state $st$ is a list of transaction symbols, ordered first partially as follows: $\mathcal{N}\preceq{}\mathcal{E}\preceq{}\mathcal{F}\preceq{}\{\mathcal{P},\mathcal{L},\mathcal{C}\}$. $\mathcal{E}$ refers to an empty slot. State slots of the same symbols are independently instantiated into transactions, except for $\mathcal{P},\mathcal{C},\mathcal{L}$. Child transactions (i.e., $\mathcal{C},\mathcal{L}$) are appended to their parent transaction of the same sender (i.e., $\mathcal{P}$). Across different senders, symbols are ordered by the parent’s GasPrice. For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolic state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
@@
In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolic end state $st^{\prime}$. Positive feedback on state $st^{\prime}$ is determined conjunctively by two metrics: 1) State coverage that indicates state $st^{\prime}$ is not covered in the corpus, and 2) state promising-ness that indicates how promising the current state is to reach a state satisfying bug oracles.
$feedback(st^{\prime},st,sdb)=1$ , if $\text{if }st\_coverage(st^{\prime},sdb)=1\land{}$ (3)
$st\_promising(st^{\prime},st)=1$
@@
Specifically, state coverage is determined by straightforwardly comparing symbolic state $st^{\prime}$, as an ordered list of symbols, with all the symbolic states in the corpus. For instance, state $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$ is different from $\mathcal{F}\mathcal{P}\mathcal{P}\mathcal{C}$. This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolic state.
@@
How promising a state $st^{\prime}$ is (i.e., $st\_promising(st^{\prime})$) is determined as follows: A mutated state $st^{\prime}$ is more promising than an unmutated state $st$ if any one of the three conditions is met: 1) State $st^{\prime}$ stores fewer normal transactions (under symbol $\mathcal{N}$) than state $st$, implying more transactions evicted and higher damage (i.e., $evict\_normal(st^{\prime},st)=1$). 2) State $st^{\prime}$ declines (speculatively) more incoming normal transactions than state $st$, also implying higher damage (i.e., $decline\_normal(st^{\prime},st)=1$). 3) The total fees of adversarial transactions in state $st^{\prime}$ are lower than those in state $st$, implying lower attack costs. We consider two forms of costs: concrete cost and symbolic cost. The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolic state $st^{\prime}$. The latter, denoted by $opcost(st^{\prime})$, optimistically estimates the cost of transactions in the current state $st^{\prime}$ that contributes to a future state triggering test oracle. We will describe the symbolic cost next. Formally, state promising-ness is calculated by Equation 4.
$st\_promising(st^{\prime},st)=$ $1,$ if $\text{if }evict\_normal(st^{\prime},st)=1\lor{}$ (4)
$decline\_normal(st^{\prime},st)=1\lor{}$
@@
Second, we use the symbolic cost to estimate how promising a state is and use it in seed energy.

### v3

@@ 4.3 Bug Oracle: Definitions & Rationale
A transaction admission timeline, $\langle{}st_{0},dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, is a successful ADAMS eviction attack, if.f. 1) the admission timeline causes full damage, that is, initial state $st_{0}$ contains only benign transactions, $ops$ are adversarial transactions, and the result end state $st_{n}$ contains only adversarial transactions (in Equation 1), and 2) the attack cost is asymmetrically low, that is, the total adversarial transaction fees in the end state $st_{n}$ to be charged are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the fees of evicted transactions in the initial state $st_{0}$ (in Equations 2 and 3). Formally,
$st_{0}\cap{}st_{n}$ $=$ $\varnothing$ (1)
$asym_{E}(st_{0},ops)$ ]$\stackrel{{\scriptstyle[}}{{]}}{def}{=}$ $\frac{\sum_{tx\in{}st_{n}}tx.fee}{\sum_{tx\in{}st_{0}}tx.fee}$ (2)
$asym_{E}(st_{0},ops)$ $<$ $\epsilon$ (3)
@@
A transaction admission timeline, $\langle{}st_{0}=\varnothing,dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, is a successful ADAMS locking attack, if.f. 1) the admission timeline causes full damage, that is, the mempool is first occupied by adversarial transactions (i.e., $st_{n}$ contains only adversarial transactions) and then declines the arriving normal transactions (i.e., $dc_{n}$ contains only normal transactions); the normal transactions and adversarial transactions are sent by different accounts (i.e., Equation 4), and 2) the attack cost is asymmetrically low, that is, the average adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the average fees of victim transactions declined (i.e., Equations 5 and 6). Formally,
$\cup_{tx\in{}dc_{n}}{tx.sender}$$\cap$ $\cup_{tx\in{}st_{n}}{tx.sender}=\varnothing$ (4)
$asym_{D}(ops)$ ]$\stackrel{{\scriptstyle[}}{{]}}{def}{=}$ $\frac{\sum_{tx\in{}st_{n}}tx.fee/\|st_{n}\|}{\sum_{tx\in{}dc_{n}}tx.fee/\|dc_{n}\|}$ (5)
$asym_{D}(ops)$ $<$ $\lambda$ (6)
Design rationale: Both definitions use the normal transaction fees to measure the attack damage because these fees are not collected by the validator in the timeline under attack and are collectible without attack. Specifically, the damage in Definition 4.2 is measured by the fees of normal transactions in $st_{0}$ which are evicted in end state $st_{n}$ and which would not have been evicted had there been no adversarial transactions in $ops$. The damage in Definition 4.3 is measured by the fees of normal transactions in $ops$ which are declined from end state $st_{n}$ and which would not have been declined had there been no adversarial transactions in $ops$.
Both definitions are strict and require causing full damage, such as evicting or declining all normal transactions. Strict definitions are necessary to ensure that what is found as exploits on the small MUT is true positive and works on the actual mempool.
The targeted attacks by this work are denial of mempool service in feeding downstream validators with valid transactions sent from benign users. They are different from other forms of DoS in blockchains, including resource exhaustion via executing malicious smart contracts [35, 27].
Particularly, in the threat model of mempool DoS, we assume the validators are benign and functional in building blocks. Because in all Ethereum clients we know, transactions are admitted into mempool without execution (i.e., executing the smart contracts they invoke), GasUsed is not a factor in transaction admission. [^2] The further design rationale for admitting transactions without execution is that an Ethereum transaction $tx$’s GasUsed is non-deterministic unless the ordering of $tx$ w.r.t. other transactions in the block is fixed and $tx$ is executed. Thus, our bug oracle does not vary GasUsed. Instead, all the transactions are fixed at $21000$ Gas (i.e., as if they don’t invoke smart contracts). [^3] We may use price and fees interchangeably. This design significantly reduces the search space without loss in covering different mempool behaviors. Besides, our bug oracle does not model the effect of EIP1559 or $BasePrice$. Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $PriorityPrice$ or damage on validator revenue.
@@ 5 Stateful Mempool Fuzzing by mpfuzz
5.1 Transaction Symbolization
The key challenge in designing mpfuzz is the large input space of transaction sequences. Recall that an Ethereum transaction consists of multiple attributes (sender, nonce, GasPrice, value, etc.), each defined in a large domain (e.g., 64-bit string). Exploring the raw transaction space is hard; the comprehensive search is inefficient, and randomly trying transactions as done in the state-of-the-art blockchain fuzzers (in § 2) is ineffective.
Intuition: We propose symbolizing transactions for efficient and effective mempool stateful fuzzing. Our key idea is to map each group of concrete transactions, triggering an equivalent mempool behavior into a distinct symbol so that searching for one transaction is sufficient to cover all other transactions under the same symbol. Transaction symbolization is expected to reduce the search space from possible concrete transactions to the symbol space.
In this work, we manually design seven symbols to represent transactions based on the Ethereum “semantics”, that is, how different transactions are admitted by the mempool.
Table 2: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
Symbol Description Tx $cost()$ $opcost()$
$\mathcal{N}$ Benign $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$ $3$ $3$
$\mathcal{F}$ Future $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$ 00 00
$\mathcal{P}$ Parent $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m+3]}\end{bmatrix}$}}$ $[4,m+3]$ $[4,m+3]$
$\mathcal{C}$ Child $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ $m+4$ $1$
$\mathcal{O}$ Overdraft $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ 00 00
$\mathcal{L}$ Latent overdraft $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ 00 00
$\mathcal{R}$ Replacement $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+4}\end{bmatrix}$}}$ $m+4$ 00
Specification: Suppose the current state contains adversarial transactions sent from $r$ accounts $A_{1},\dots,A_{r}$. These accounts have an initial balance of $m$ Ether, where $m$ equals the mempool capacity (say $m=1000$). The rationale is that these accounts can send at most $m$ adversarial transactions, each minimally spending one Ether, to just occupy a mempool of $m$ slots. A larger value of attacker balance is possible but increases the search space.
@@ 5.1 Transaction Symbolization
Symbol $\mathcal{F}$ defines the transaction subspace of any future transaction sent from an adversarial account; the future transaction is defined w.r.t. the current state. mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, GasPrice $m+4$ wei, and any adversarial account $A*$. The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$.
Symbol $\mathcal{P}$ defines the transaction subspace of any parent transaction from an adversarial account w.r.t. the current state. mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m+3]}\end{bmatrix}$}}$. Here, the transaction is fixed with a GasPrice in the range $[4,m+3]$ wei, value at $1$ Ether, and nonce at $1$. The GasPrice of a future transaction ($\mathcal{F}$) is $m+4$ wei, which is higher than the price of a parent transaction ($\mathcal{P}$) in $[4,m+3]$ wei. The purpose is to ensure that a future transaction can evict any parent transaction during fuzzing.
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$. The instantiated child transaction is fixed at GasPrice of $m+4$
Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$. Recall that all adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{L}$ defines the transaction subspace of any adversarial latent-overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$. All adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{R}$ defines the transaction subspace of any adversarial “replacement” transaction w.r.t., the state. Unlike the symbols above, the transactions under Symbol $\mathcal{R}$ must share the same sender and nonce with an existing transaction in the current state $st$. mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+4}\end{bmatrix}$}}$. Here, we simplify the problem and only consider replacing the transaction with nonce $1$.
@@
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 2, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$). The design rationale of transaction symbolization is deferred to Appendix A.1.
@@ 5.1.1 State Search Algorithm
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [22]). A seed consists of a symbolized input $ops$, which consists of transaction symbols, and a symbolized state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
@@
Mutation feedback: We describe how mempool states are symbolized before presenting state-based feedback. In mpfuzz, a symbolized state $st$ is a list of transaction symbols, ordered first partially as follows: $\mathcal{N}\preceq{}\mathcal{E}\preceq{}\mathcal{F}\preceq{}\{\mathcal{P},\mathcal{L},\mathcal{C}\}$. $\mathcal{E}$ refers to an empty slot. State slots of the same symbols are independently instantiated into transactions, except for $\mathcal{P},\mathcal{C},\mathcal{L}$. Child transactions (i.e., $\mathcal{C},\mathcal{L}$) are appended to their parent transaction of the same sender (i.e., $\mathcal{P}$). Across different senders, symbols are ordered by the parent’s GasPrice. For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolized state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
@@
In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolized end state $st^{\prime}$. Positive feedback on state $st^{\prime}$ is determined conjunctively by two metrics: 1) State coverage that indicates state $st^{\prime}$ is not covered in the corpus, and 2) state promising-ness that indicates how promising the current state is to reach a state satisfying bug oracles.
$feedback(st^{\prime},st,sdb)$ $=$ $st\_coverage(st^{\prime},sdb)==1\land{}$ (7)
$st\_promising(st^{\prime},st)==1$
@@
Specifically, state coverage is determined by straightforwardly comparing symbolized state $st^{\prime}$, as an ordered list of symbols, with all the symbolized states in the corpus. For instance, state $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$ is different from $\mathcal{F}\mathcal{P}\mathcal{P}\mathcal{C}$. This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolized state.
@@
How promising a state $st^{\prime}$ is (i.e., $st\_promising(st^{\prime})$) is determined as follows: A mutated state $st^{\prime}$ is more promising than an unmutated state $st$ if any one of the three conditions is met: 1) State $st^{\prime}$ stores fewer normal transactions (under symbol $\mathcal{N}$) than state $st$, implying more transactions evicted and higher damage (i.e., $evict\_normal(st^{\prime},st)=1$). 2) State $st^{\prime}$ declines (speculatively) more incoming normal transactions than state $st$, also implying higher damage (i.e., $decline\_normal(st^{\prime},st)=1$). 3) The total fees of adversarial transactions in state $st^{\prime}$ are lower than those in state $st$, implying lower attack costs. We consider two forms of costs: concrete cost and symbolized cost. The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolized state $st^{\prime}$. The latter, denoted by $opcost(st^{\prime})$, optimistically estimates the cost of transactions in the current state $st^{\prime}$ that contributes to a future state triggering test oracle. We will describe the symbolized cost next. Formally, state promising-ness is calculated by Equation 8.
$st\_promising(st^{\prime},st)$ $=$ $evict\_normal(st^{\prime},st)==1\lor{}$ (8)
$decline\_normal(st^{\prime},st)==1\lor{}$
@@
Second, we use the symbolized cost to estimate how promising a state is and use it in seed energy.